OpenStudy (anonymous):

use partila derivatis to find dy/dx when y is given implicitly by the equation x^2y-10x=sqrt(xy) + y

OpenStudy (anonymous):

\[x^2y-10x=\sqrt{xy}+y\\ {d\over dx}(x^2y)-10{d\over dx}(x)={d\over dx}\sqrt{xy}+{dy\over dx} \] use product rule and chain rules for the two of them

OpenStudy (anonymous):

lets approach it step by step what is \[{d\over dx}(x^2y)=?\]

OpenStudy (anonymous):

it is a product.. so you have to use the product rule first \[{d\over dx}(x^2y)={d\over dx}(x^2)\times y+x^2\times{d\over dx}(y)\]

OpenStudy (anonymous):

2xy-10=y/2sqrt(xy) partial of x right?

OpenStudy (anonymous):

do not jump.. :) lets take it slow

OpenStudy (anonymous):

did you find the above derivative piece?

OpenStudy (anonymous):

2xy+x^2 *d/dx(y)

OpenStudy (anonymous):

great. now, how about \[{d\over dx}\sqrt{xy}=?\]

OpenStudy (anonymous):

y/sqrt(xy)

OpenStudy (anonymous):

note.. chain rule first, then product rule

OpenStudy (anonymous):

y/2sqrt(xy)

OpenStudy (anonymous):

there are two ways to do this.. let me go over the first way \[\Large{{d\over dx}\sqrt{xy}={d\over dx}(xy)^{1\over2}={1\over2}(xy)^{-1\over2}{d\over dx}(xy)\\ \quad={1\over2(xy)^{1\over2}}\left(1\times y+x\times{dy\over dx}\right)\\ \quad={1\over2\sqrt{xy}}\left(1\times y+x\times{dy\over dx}\right)\\ \quad={y\over2\sqrt{xy}}+{x\over2\sqrt{xy}}{dy\over dx}\\ \quad=\color{red}{{1\over2}\sqrt{y\over x}+{1\over2}\sqrt{x\over y}{dy\over dx}} }\]

OpenStudy (anonymous):

okay. is see the math here.

OpenStudy (anonymous):

now we plug em up in the first one. and solve for \({dy\over dx}\)

OpenStudy (anonymous):

before one starts should we set equation to = 0?